Alpha Company used the periodic inventory system for purchase & sales of merchandise. Discount terms for both purchase & sales are, FOB Destination, 2/10, n30 and the gross method is used.

Alpha Company purchased on account $2,500 of merchandise from Bravo Company on May 2, 2016.

Alpha Company returned, to Bravo Company, $250 of this merchandise on May 3, 2016.

Freight charges related to this transaction of $150 were paid by Bravo Company.

Use this information to prepare Alpha Company's General Journal entry (without explanation) for the payment for merchandise on May 12. If no entry is required then write "No Entry Required."
Answer #1

Cost of paochases - 2,500 pcorchase verons 250 Net puacha sos = 2,500 -250 - 2,250 Descocent teams = 4/10 ner 30, which means